# Robust Deep Reinforcement Learning against Adversarial Behavior Manipulation

### v2

@@ 1 Introduction
Applications of Deep Reinforcement Learning (DRL) have grown significantly in recent years [1, 2]. However, when DRL is deployed in mission-critical tasks [3, 4, 5], it is crucial to understand its susceptibilities to adversarial attacks and deal with them. One such vulnerability is the manipulation of a victim’s behavior by an adversary that can deceive the state observed by a victim agent [6, 7, 8].
The primary objective of this study is to introduce behavior-targeted attacks and corresponding countermeasures. In this attack, the adversary’s goal is to steer the victim’s trained policy toward a target policy specified by the adversary. Behavior-targeted attacks introduce novel threats that cannot be realized by conventional reward-minimization attacks, where the adversary’s goal is to minimize the victim’s reward [9, 10, 11, 12]. Moreover, as we detail later, existing defenses tailored to counter reward-minimization attacks are ineffective against these novel threats, which underscores the necessity of dedicated countermeasures. Below, we illustrate two scenarios where the behavior-targeted attacks are particularly relevant.
Scenario 1. Consider an autonomous vehicle controlled by a DRL agent that receives higher rewards for reaching its destination quickly and safely, but incurs a large penalty if it causes an accident [3, 13]. Reward-minimization attacks would aim to induce an accident at some point along the vehicle’s route. In contrast, behavior-targeted attacks can take various forms, regardless of the victim’s reward. For instance, the adversary could manipulate the vehicle to slow down at a specific crossing point, creating congestion in a strategically chosen region. Alternatively, the adversary might steer the autonomous vehicle toward a particular store to generate economic benefit for the adversary.
Scenario 2. Consider a recommendation system powered by DRL [14, 15] that maximizes user satisfaction (reward) by observing the user’s purchase history and recommending relevant items. In this context, reward-minimization attacks attempt to degrade user satisfaction by generating suboptimal recommendations. In contrast, behavior-targeted attacks manipulate the system to serve specific objectives, such as prompting certain products that benefit the adversary or suppressing the recommendation of items specified by the adversary.
@@
Although several behavior-targeted attacks have been proposed [16, 17, 18, 19], they all share a critical limitation: each requires white-box access to the victim’s policy, including its architecture and parameters. In many real-world scenarios, the adversary is unlikely to have full knowledge of the victim’s policy, making this assumption severely limit practical applications.
@@
Building on this formulation, we propose the Behavior Imitation Attack (BIA), a novel attack framework that is applicable even when access to the victim’s policy is severely limited. Since the reformulated objective can be optimized via demonstration-based imitation learning, we utilize established algorithms [20, 21, 22]. The advantage of leveraging imitation learning is that BIA enables training adversarial policies directly from demonstrations of the desired behavior, eliminating the need for reward modeling. The adversary can readily prepare demonstrations necessary for BIA simply by performing the desired behavior several times. We empirically show that even under the most restrictive no-box setting, where the adversary cannot observe any output of the victim’s policy, BIA achieves attack performance competitive with baselines requiring white-box access.
@@
Most existing defense strategies, such as adversarial training [11, 23, 8, 24], certified defenses [25, 26, 27, 28, 29], and regret-based robust learning [30, 31, 32], are not readily adaptable to counter behavior-targeted attacks. This is because the adversary’s target policy is unknown to the defender and determined independently of the victim’s reward. Although policy smoothing [33, 7] may offer some degree of robustness by stabilizing the policy’s outputs against adversarial perturbations, their theoretical robustness guarantees are limited to reward-minimization attacks and not directly extended to behavior-targeted attacks. Moreover, policy smoothing often degrades performance on the victim’s original tasks, as it imposes excessive constraints on the policy’s representational capacity.
@@
Motivated by these insights, we introduce Time Discounting Robust Training (TDRT), which incorporates a time-discounted regularization term to suppress the sensitivity of the policy’s actions. This regularization more strongly reduces the sensitivity at critical earlier stages of trajectories and progressively weakens at later stages. By concentrating regularization in early trajectories, TDRT maintains the policy’s representational capacity and thus preserves original-task performance, while simultaneously enhancing robustness against behavior-targeted attacks. In our experiments, compared to uniform policy-smoothing defenses without time-discounting, TDRT improves original-task performance by 28.2% while maintaining comparable robustness. To the best of our knowledge, TDRT is the first defense specifically designed for behavior-targeted attacks.
@@
  • We demonstrate the effectiveness of our proposed attack and defense methods using the Meta-World [34], MuJoCo [35], and MiniGrid [36] environments.
@@ 2 Related works
Attack methods.
Reward-minimization attacks aim to minimize the victim’s cumulative reward and have been extensively studied [6, 9, 10, 11, 8, 12]. Some approaches [37, 38] steer the victim’s policy toward a specific policy which minimizes the cumulative reward via Projected Gradient Descent (PGD) [39], but these require white-box access. As another direction, enchanting attacks seek to lure the victim to a predetermined terminal state without prescribing its entire trajectory [40, 41, 42]. Some approaches [40, 41] optimize per-step perturbations to guide the victim to the goal, and Ying et al. [42] introduce a universal-perturbation approach. Neither reward-minimization attacks nor enchanting attacks are designed to manipulate the victim’s entire behavior. Although some works focus on behavior-targeted attacks [16, 17, 18, 19], they all require white-box access to the victim’s policy, as mentioned above. Moreover, attacks such as those in [16, 17], which were designed for specific domains (e.g., autonomous driving), have limited generalizability. Bai et al. [18, 19] proposed a non-heuristic attack based on preference-based imitation learning. However, it still relies on white-box access and involves three separate stochastic optimization processes, making it highly resource-intensive. Several studies have proposed poisoning attacks that intervene during the victim’s training phase [43, 44, 45, 46, 47]. However, they differ from our attacks on the trained and fixed victims.
Defense methods.
@@
Adversarial training [11, 23, 8, 24] involves training the agent to be robust against adversarial perturbations generated by a hypothetical adversary. However, its effectiveness against behavior-targeted attacks is limited, as the defender cannot anticipate which behavior the adversary aims to impose. Existing policy smoothing methods [33, 7] are designed for reward-minimization attacks and significantly degrade performance on the original tasks as previously noted. Other approaches include regret-based robust learning [30, 31, 32] and certified defenses [25, 26, 27, 28, 29]. Regret is defined as the difference between the victim’s attacked and unattacked rewards, and regret-based robust learning focuses on minimizing this difference. Certified defenses guarantee a lower bound on the victim’s reward under attack. As they both use reward-based metrics and are designed for reward-minimization attacks, they are less effective against behavior-targeted attacks that are independent of the victim’s reward. Liu et al. [48] proposed an adaptive defense method robust against multiple types of attacks, not limited to reward-minimization attacks. In their method, the defender prepares multiple robust policies and then selects the policy that maximizes the victim’s reward under attack based on rewards obtained in previous episodes. However, our approach focuses on a single static victim in a stationary environment, so their adaptive setting differs from ours. For a more detailed and comprehensive review of related work, please refer to Appendix A.
@@ 3 Preliminaries
To model the situation where an adversary intervenes in the victim’s state observations, we use SA-MDP [7]. Let the victim follow a fixed policy $\pi$ in the MDP $M$. In an SA-MDP, the adversary introduces an adversarial policy $\nu:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ that interferes with the victim’s state observations by making the victim observe a false state $\hat{s}\sim\nu(\cdot|s)$ at each time step without altering the true state of the environment (see Figure 1).
@@ 4 Threat model
We assume the adversary does not have full access to the victim’s policy $\pi$, including its neural network parameters and training algorithm. Specifically, we define two access models: black-box and no-box. In the black-box setting, the adversary can only observe the victim’s policy inputs (i.e., observed states) and the corresponding outputs (i.e., actions taken) at each time step. In the no-box setting, the adversary observes only the policy inputs at each time step. Since the no-box setting provides less information than the black-box setting, it presents a greater challenge to the adversary.
@@
We assume that the adversary can alter the victim’s state observation from the true state $s$ to the false state $\hat{s}\in\mathcal{B}(s)$. This assumption is common in most adversarial attacks on RL [7, 11, 8].
@@ 5 Attack method
First, we define a composite policy of the victim’s policy and the adversarial policy to define the adversarial objective. Let $s\in\mathcal{S}$ be a true state. The adversarial policy provides the falsified observation $\hat{s}\sim\nu(\cdot|s)$, and then the victim selects the next action by $a\sim\pi(\cdot|\hat{s})$. Consequently, the victim’s resulting behavior policy under the adversary’s influence is represented by the composite policy ${\pi\circ\nu}$:
${\pi\circ\nu}(a|s)\triangleq\sum_{\hat{s}\in\mathcal{S}}\nu(\hat{s}|s)\pi(a|\hat{s}).$ (1)
@@
where $\mathcal{D}$ is some divergence measure between two policies. Unfortunately, directly optimizing Equation (2) is infeasible when the adversary does not have white-box access to the victim policy $\pi$, since the gradients of ${\pi\circ\nu}$ cannot be computed.
To solve Equation (2), we provide a novel theoretical result that, under a mild assumption, Equation (2) can be reformulated into an problem that maximizes the cumulative reward in another MDP $\hat{M}$ in which the adversarial policy itself serves as the policy to be optimized:
@@
where $f$ and $g$ are arbitrary convex and concave functions, respectively, and $d:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a discriminator. Let $d_{\star}$ be the optimal discriminator in Equation (3). Under this assumption, define the reward function $\hat{R}_{d}$ and the state transition probability $\hat{p}$ as follows:
@@
The proof is inspired by Lemma 1 of [7] and is provided in Appendix B.1. Crucially, any standard RL algorithm can maximize the cumulative reward under MDP $\hat{M}$ without requiring white-box access to $\pi$ since the policy for MDP $\hat{M}$ depends solely on $\nu$ rather than on the composite ${\pi\circ\nu}$. The dynamics of the MDP $\hat{M}$ encapsulate the victim’s policy, and this idea has been considered in [49, 19]. However, their idea were not theoretically justified because [49] only introduces the concept abstractly without empirical validation or theoretical analysis, and [19] lacks a proof and concrete reward design. In this study, we provide a complete proof and present a concrete reward design for learning the optimal adversarial policy for the first time.
@@
Building on the above insight, we propose Behavior Imitation Attack (BIA). Specifically, we bridge the behavior-targeted attacks and established imitation learning [20, 50, 22] via Theorem 5.1. We then present a practical algorithm for implementing BIA.
Imitation learning (IL) trains a policy that mimics an expert policy $\pi_{E}$ from given demonstrations without requiring rewards from the environment. Its objective can be expressed analogously Equation (2):
@@
By using IL algorithms that satisfy the assumptions of Theorem 5.1, we can carry out attacks even under black-box and no-box settings. Specifically, many IL algorithms rely on divergences that admit a variational representation. Thus, by reformulating the adversary’s objective using Theorem 5.1, we can leverage these well-established methods to train the adversarial policy.
As a concrete example of an IL algorithm that holds the assumption, we illustrate how to apply BIA to GAIL[20], a representative ILfD method. In GAIL, the adversary’s objective function can be rewritten in the following variational form using a discriminator $D\colon\mathcal{S}\times\mathcal{A}\to[0,1]$:
@@
The second term is computed with demonstration $\tau_{\text{tgt}}$ instead of $\pi_{\text{tgt}}$ in practice. In Appendix C, we show that under our problem setting, the adversary’s objective function can be reduced to a distribution matching problem. Consequently, we confirm that Equation (8) holds exactly. By Theorem 5.1, this then reduces to the following optimization problem:
@@
where $\hat{R}_{D}$ is the function obtained by replacing the discriminator $g(d_{\star})$ with $\log D_{\star}$ in Equation (4). This is a standard RL problem and can be solved using any RL algorithm. In Appendix D, we further show how BIA extends to other approaches, including ILfO [50] and the state-of-the-art ILfD algorithm [22].
@@
We present a practical algorithm for BIA that utilizes GAIL to train an adversarial policy in Algorithm 1. Since $\hat{R}$ in Equation (4) represents the conditional expectation of $-g(d_{\star}(s,a,s^{\prime}))$ for the adversarial policy, we simply approximate it by a single sample $g(d_{\star}(s,a,s^{\prime}))=\log D(s,a)$ at each step. This approximation preserves exactness in expectation, so there is no discrepancy between Theorem 5.1 and Algorithm 1. As our approach does not require direct access to the victim’s policy, BIA can be applied in a black-box setting. Furthermore, when ILfO is used, the discriminator $D$ only needs states and the next states to train the adversarial policy. Since it does not require the victim’s policy’s selected actions, our method is applicable even in a no-box setting. Also, to stabilize the learning process, we avoid imposing large negative rewards; instead, we limit the range of false states, ensuring that the adversarial policy is enforced to select falsified states in $\mathcal{B}(s)$ in our experiments.

### v3

@@ 1 Introduction
Applications of Deep Reinforcement Learning (DRL) have grown significantly in recent years (Berner et al., 2019; Ouyang et al., 2022; Guo et al., 2025). However, when DRL is deployed in mission-critical tasks (Sallab et al., 2017; Degrave et al., 2022; Yu et al., 2021), it is crucial to understand its susceptibilities to adversarial attacks and deal with them. One such vulnerability is the manipulation of a victim’s behavior by an adversary that can deceive the state observed by a victim agent (Huang et al., 2017; Zhang et al., 2020b; Sun et al., 2022).
The primary objective of this study is to introduce behavior-targeted attacks and corresponding countermeasures. In this attack, the adversary’s goal is to steer the victim’s trained policy toward a target policy specified by the adversary. Behavior-targeted attacks introduce novel threats that cannot be realized by conventional reward-minimization attacks, where the adversary’s goal is to minimize the victim’s reward (Pattanaik et al., 2018; Zhang et al., 2021; McMahan et al., 2024). Moreover, as we detail later, existing defenses tailored to counter reward-minimization attacks are ineffective against these novel threats, which underscores the necessity of dedicated countermeasures. Below, we illustrate two scenarios where the behavior-targeted attacks are particularly relevant.
Scenario 1. Consider an autonomous vehicle controlled by a DRL agent that receives higher rewards for reaching its destination quickly and safely, but incurs a large penalty if it causes an accident (Kiran et al., 2021). Reward-minimization attacks would aim to induce an accident at some point along the vehicle’s route. In contrast, behavior-targeted attacks can take various forms, regardless of the victim’s reward. For instance, the adversary could manipulate the vehicle to slow down at a specific crossing point, creating congestion in a strategically chosen region. Alternatively, the adversary might steer the autonomous vehicle toward a particular store to generate economic benefit for the adversary.
Scenario 2. Consider a recommendation system powered by DRL (Chen et al., 2023; Fu et al., 2022) that maximizes user satisfaction (reward) by observing the user’s purchase history and recommending relevant items. In this context, reward-minimization attacks attempt to degrade user satisfaction by generating suboptimal recommendations. In contrast, behavior-targeted attacks manipulate the system to serve specific objectives, such as prompting certain products that benefit the adversary or suppressing the recommendation of items specified by the adversary.
@@
Although several behavior-targeted attacks have been proposed (Hussenot et al., 2020; Boloor et al., 2020; Bai et al., 2024; 2025), they all share a critical limitation: requiring white-box access to the victim’s policy, including its architecture and parameters. In many real-world scenarios, the adversary is unlikely to have full knowledge of the victim’s policy, which severely limits practical applicability.
@@
Building on this formulation, we propose the Behavior Imitation Attack (BIA), a novel attack framework that is applicable even when access to the victim’s policy is severely limited. Since the reformulated objective can be optimized via demonstration-based imitation learning, we utilize established algorithms (Ho and Ermon, 2016; Kostrikov et al., 2018; Chang et al., 2024). The advantage of leveraging imitation learning is that BIA enables training adversarial policies directly from demonstrations of the desired behavior, eliminating the need for reward modeling. The adversary can readily prepare demonstrations necessary for BIA simply by performing the desired behavior several times. We empirically show that even under the most restrictive no-box setting, where the adversary cannot observe any output of the victim’s policy, BIA achieves attack performance competitive with baselines requiring white-box access.
@@
Most existing defense strategies, such as adversarial training (Zhang et al., 2021; Oikarinen et al., 2021; Sun et al., 2022; Liang et al., 2022), certified defenses (Wu et al., 2022; Kumar et al., 2022; Mu et al., 2024; Sun et al., 2024; Wang et al., 2025), and regret-based robust learning (Jin et al., 2018; Rigter et al., 2021; Belaire et al., 2024), are not readily adaptable to counter behavior-targeted attacks. This is because the adversary’s target policy is unknown to the defender and determined independently of the victim’s reward. Although policy smoothing (Shen et al., 2020; Zhang et al., 2020b) may offer some degree of robustness by stabilizing the policy’s outputs against adversarial perturbations, their theoretical robustness guarantees are limited to reward-minimization attacks and not directly extended to behavior-targeted attacks. Moreover, policy smoothing often degrades performance on the victim’s original tasks, as it imposes excessive constraints on the policy’s representational capacity.
@@
Motivated by these insights, we introduce Time-Discounted Robust Training (TDRT), which incorporates a time-discounted regularization term to suppress the sensitivity of the policy’s actions. This regularization more strongly reduces the sensitivity at critical earlier stages of trajectories and progressively weakens at later stages. By concentrating regularization in early trajectories, TDRT maintains the policy’s representational capacity and thus preserves original-task performance, while simultaneously enhancing robustness against behavior-targeted attacks. In our experiments, compared to uniform policy-smoothing defenses without time-discounting, TDRT improves original-task performance by 28.2% while maintaining comparable robustness. To the best of our knowledge, TDRT is the first defense specifically designed for behavior-targeted attacks.
@@
  • We demonstrate the effectiveness of our proposed attack and defense methods using the Meta-World (Yu et al., 2020), MuJoCo (Todorov et al., 2012), and MiniGrid (Chevalier-Boisvert et al., 2018) environments.
@@ 2 Related works
2.1 Attack methods
Existing attack methods on DRL agents can be broadly classified into three categories: reward-minimization attacks, enchanting attacks, and behavior-targeted attacks. Reward-minimization attacks aim to reduce the victim’s cumulative reward and have been extensively studied (Zhang et al., 2021; Sun et al., 2022; McMahan et al., 2024). As another direction, enchanting attacks aim to lure the victim into a predetermined terminal state without specifying the full trajectory (Ying et al., 2023). Since neither reward-minimization nor enchanting attacks are intended to manipulate the victim’s entire behavior, we treat them as out of scope.
Behavior-targeted attacks instead aim to steer the victim toward an adversary-specified target policy. However, existing approaches typically assume white-box access to the victim’s policy, as discussed above. In addition, some attacks (Hussenot et al., 2020; Boloor et al., 2020) are tailored to specific domains such as autonomous driving, which limits their generalizability. More recently, Bai et al. (2024; 2025) proposed a non-heuristic attack based on preference-based reinforcement learning, but it still relies on white-box access and requires three separate stochastic optimization procedures, making it highly resource-intensive.
While several studies have proposed poisoning attacks that intervene during the victim’s training phase (Sun et al., 2021; Rangi et al., 2022; Xu et al., 2023; Xu and Singh, 2023; Rathbun et al., 2024), they fall outside our threat model (see Appendix A.1 for details).
2.2 Defense methods
Adversarial training (Zhang et al., 2021; Oikarinen et al., 2021; Sun et al., 2022; Liang et al., 2022; Li et al., 2024a) optimizes the agent to be robust against adversarial perturbations generated by a hypothetical adversary. However, the effectiveness of adversarial training against behavior-targeted attacks is limited, as the defender cannot anticipate which behavior the adversary aims to impose. Existing policy smoothing methods (Shen et al., 2020; Zhang et al., 2020b) are designed for reward-minimization attacks and significantly degrade performance on the original tasks as previously noted.
Other approaches include regret-based robust learning (Jin et al., 2018; Rigter et al., 2021; Belaire et al., 2024) and certified defenses (Wu et al., 2022; Kumar et al., 2022; Mu et al., 2024; Sun et al., 2024; Wang et al., 2025). Regret is defined as the difference between the victim’s attacked and unattacked rewards, and regret-based robust learning focuses on minimizing this difference. Certified defenses guarantee a lower bound on the victim’s reward under attack. As they both use reward-based metrics and are designed for reward-minimization attacks, they are less effective against behavior-targeted attacks that are independent of the victim’s reward.
@@ 2.2 Defense methods
Liu et al. (2024) proposed an adaptive defense method robust against multiple types of attacks, not limited to reward-minimization attacks. In their method, the defender prepares multiple robust policies and then selects the policy that maximizes the victim’s reward under attack based on rewards obtained in previous episodes. However, our approach focuses on a single static victim in a stationary environment, so their adaptive setting differs from ours.
Several works (Sun and Zheng, 2024; YANG and Xu, 2024) propose diffusion-based defenses, which perform purification at inference time to remove adversarial perturbations from corrupted states. In contrast, our approach is a training-time regularization method and requires no additional preprocessing during inference. As a result, these defenses operate on different components of the pipeline, making them orthogonal and complementary. For a more detailed and comprehensive review of related work, please refer to Appendix A.
@@ 3 Preliminaries
To model the situation where an adversary intervenes in the victim’s state observations, we use SA-MDP (Zhang et al., 2020b). Let the victim follow a fixed policy $\pi$ in the MDP $M$. In an SA-MDP, the adversary introduces an adversarial policy $\nu:\mathcal{S}\rightarrow\mathcal{P}(\mathcal{S})$ that interferes with the victim’s state observations by making the victim observe a false state $\hat{s}\sim\nu(\cdot|s)$ at each time step without altering the true state of the environment (see Figure 1).
@@ 4 Threat model
We assume the adversary does not have full access to the victim’s policy $\pi$, including its neural network parameters and training algorithm. Specifically, we define two access models: black-box and no-box. In the black-box, the adversary can only observe the victim’s policy inputs (i.e., observed states) and the corresponding outputs (i.e., actions taken) at each time step. In the no-box, the adversary observes only the policy inputs at each time step. Since the no-box setting provides less information than the black-box, it presents a greater challenge to the adversary.
@@
We assume that the adversary can alter the victim’s state observation from the true state $s$ to the false state $\hat{s}\in\mathcal{B}(s)$. This assumption is common in most adversarial attacks on RL (Zhang et al., 2020b; 2021; Sun et al., 2022).
@@ 5 Attack method
[…]
@@
where $\mathcal{D}$ is some divergence measure between two policies.
Unfortunately, existing imitation learning algorithms are not directly applicable for optimizing equation 2 under our threat model. Typical imitation learning methods treat the composite policy $\pi_{\nu}$ as the optimizable policy and assume it is directly updatable via gradients. However, since the adversary lacks white-box access to the victim policy $\pi$, it cannot backpropagate through $\pi$ to compute the gradient of the objective with respect to $\nu$. Therefore, this assumption does not hold.
To overcome this difficulty, we provide a novel theoretical result that, under a mild assumption, equation 2 can be reformulated into a problem that maximizes the cumulative reward in another MDP $\hat{M}$ in which the adversarial policy itself serves as the policy to be optimized:
@@
where $f$ and $g$ are arbitrary convex and concave functions, respectively, and $d:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ is a discriminator. Let $d_{\star}$ be the optimal discriminator in equation 3. Under this assumption, define the reward function $\hat{R}_{d}$ and the state transition probability $\hat{p}$ as follows:
@@
The proof is provided in Appendix B.1. Crucially, any standard RL algorithm can maximize the cumulative reward under MDP $\hat{M}$ without requiring white-box access to $\pi$ since the policy for MDP $\hat{M}$ depends solely on $\nu$ rather than on the composite $\pi\circ\nu$}. The dynamics of the MDP $\hat{M}$ encapsulate the victim’s policy, and this idea has been considered in (Schott et al., 2024; Bai et al., 2025). However, their idea was not theoretically justified because (Schott et al., 2024) only introduces the concept abstractly without empirical validation or theoretical analysis, and (Bai et al., 2025) lacks a proof and concrete reward design. We provide a complete proof and present a concrete reward design for learning the optimal adversarial policy for the first time.
@@
Building on the above insight, we propose Behavior Imitation Attack (BIA). Specifically, we bridge the behavior-targeted attacks and established Imitation Learning (IL) (Ho and Ermon, 2016; Torabi et al., 2019; Chang et al., 2024) via Theorem 5.1. We then present a practical algorithm for implementing BIA.
IL trains a policy that mimics an expert policy $\pi_{E}$ from given demonstrations without requiring rewards from the environment. Its objective can be expressed analogously equation 2:
@@
By using IL algorithms that satisfy the assumptions of Theorem 5.1, we can carry out attacks even under black‐box and no‐box settings. Specifically, many IL algorithms rely on divergences that admit a variational representation. Thus, by reformulating the adversary’s objective using Theorem 5.1, we can leverage these well-established methods to train the adversarial policy.
As a concrete example of an IL algorithm that holds the assumption, we illustrate how to apply BIA to GAIL(Ho and Ermon, 2016), a representative ILfD method. In GAIL, the adversary’s objective function can be rewritten in the following variational form using a discriminator $D\colon\mathcal{S}\times\mathcal{A}\to[0,1]$:
@@
The second term is computed with demonstration $\tau_{\text{tgt}}$ instead of $\pi_{\text{tgt}}$ in practice. In Appendix C, we show that under our problem setting, the adversary’s objective function can be reduced to a distribution matching problem. Consequently, we confirm that equation 8 holds exactly. By Theorem 5.1, this then reduces to the following optimization problem:
@@
where $\hat{R}_{D}$ is the function obtained by replacing the discriminator $g(d_{\star})$ with $\log D_{\star}$ in equation 4. This is a standard RL problem and can be solved using any RL algorithm. In Appendix D, we further show how BIA extends to other approaches, including ILfO (Torabi et al., 2019) and the state-of-the-art ILfD algorithm (Chang et al., 2024).
@@
We present a practical algorithm for BIA that utilizes GAIL to train an adversarial policy in Algorithm 1. Since $\hat{R}$ in equation 4 represents the conditional expectation of $g(d_{\star}(s,a,s^{\prime}))$ for the adversarial policy, we simply approximate it by a single sample $g(d_{\star}(s,a,s^{\prime}))=\log D(s,a)$ at each step. This approximation preserves exactness in expectation, so there is no discrepancy between Theorem 5.1 and Algorithm 1. As our approach does not require direct access to the victim’s policy, BIA can be applied in a black-box setting. Furthermore, when ILfO is used, the discriminator $D$ only needs states and the next states to train the adversarial policy. Since it does not require the victim’s policy’s selected actions, our method is applicable even in a no-box setting. Also, to stabilize the learning process, we avoid imposing large negative rewards; instead, we limit the range of false states, ensuring that the adversarial policy is enforced to select falsified states in $\mathcal{B}(s)$ in our experiments.